Find the number of solutions of sin theta + 2 cos theta =0 over the interval [-pi,pi/2].

Answer by stanbon(75887)

Find the number of solutions of
sin theta + 2 cos theta =0
over the interval [-pi,pi/2].
----
Divide thru by cos(t) to get:
tan(t) + 2 = 0
tan(t) = -2
------
t = arctan(-2) = -1.1070 in QIV
-----------------
t = -1.1070+pi = 2.0344 in QII
------------------------------------
Cheers,
Stan H.
==================